# MAAT: Mamba Adaptive Anomaly Transformer with association discrepancy for time series

### v1

@@ 6.2 Reconstruction loss
where $i$ denotes the batch index, $L^{AT}_{i}$ represents the Anomaly Transformer reconstruction loss, and $L^{MAAT}_{i}$ represents our method reconstruction loss at batch $i$. The resulting differential analysis, visualized in Figure 2, employs a color-coded representation where green columns indicate regions where our method achieves lower reconstruction loss ($\Delta L_{i}>0$). In contrast, red columns denote areas where the Anomaly Transformer performs better ($\Delta L_{i}<0$). This visualization effectively demonstrates the comparative advantages of our approach across different batches of the training process.
MAAT achieves superior anomaly detection on the SMD dataset (F1: 92.30% vs. baseline 87.18%) by minimizing reconstruction loss (Figure 2(a)), reflecting its ability to model standard patterns accurately. The logarithmic difference (Figure 2(b)) shows predominant green bars ($\Delta L_{i}>0$), indicating MAAT’s consistent advantage through Sparse Attention (focusing on critical time steps) and Mamba-SSM (capturing long-range dependencies). Transient red bars ($\Delta L_{i}<0$) align with minor precision dips (89.03% vs. DCdetector 85.82%) but do not compromise overall robustness. Enhanced reconstruction fidelity reduces false positives/negatives, driving higher Affiliation Recall (93.51%) and Range-based Recall (79.11%), while gated attention ensures stability in noisy, non-stationary conditions. This synergy validates MAAT efficacy for industrial monitoring. Ability in noisy, non-stationary conditions. This synergy validates MAAT efficacy for industrial monitoring.
MAAT outperforms the Anomaly Transformer on the MSL dataset by achieving lower reconstruction loss Figure 2(a), reflecting its enhanced ability to model noisy, variable patterns from Mars rover sensors. The logarithmic difference plot ($\Delta L_{i}$, right) is dominated by green bars ($\Delta L_{i}>0$), demonstrating MAAT consistent superiority through Mamba-SSM, which captures long-range dependencies in transient telemetry data, and Sparse Attention, which reduces redundancy in short-term noisy windows. Transient red bars ($\Delta L_{i}<0$) correlate with minor precision trade-offs (92.06% vs. Anomaly Transformer 91.02%) but are offset by significantly higher recall (98.33% vs. 96.03%). This precision-recall balance, coupled with robust Affiliation Recall (96.49%) and Range-based Recall (90.97%), validates MAAT resilience to non-stationarity, critical for space mission monitoring systems.
MAAT achieves superior anomaly detection on the SMAP dataset (F1: 96.99% vs. Anomaly Transformer 96.41%) by maintaining lower reconstruction loss Figure 2(f), reflecting its ability to discern subtle anomalies in satellite telemetry data. The logarithmic difference plot ($\Delta L_{i}$, right) shows predominant green bars ($\Delta L_{i}>0$), driven by Mamba-SSM capacity to model long-range dependencies in SMAP sparse, noisy signals and Sparse Attention focus on critical temporal windows. While MAAT precision (94.75% vs. 93.59%) and F1 improve, a minor recall trade-off (99.33% vs. 99.41%) aligns with rare red bars ($\Delta L_{i}<0$), where transient anomalies mimic normal patterns. This precision-centric gain enhances Affiliation Precision (49.34%) and Volume-based ROC (92.06%), prioritizing accurate anomaly localization over exhaustive detection. MAAT robustness in noisy, long-sequence contexts validates its utility for satellite monitoring systems requiring reliable anomaly identification.
MAAT achieves superior anomaly detection on the PSM dataset (F1: 98.32% vs. Anomaly Transformer 97.47%) by maintaining lower reconstruction loss Figure 2(c), reflecting its ability to handle high-dimensional, noisy sensor data from server machines. The logarithmic difference plot ($\Delta L_{i}$, right) is dominated by green bars ($\Delta L_{i}>0$), driven by Sparse Attention focus on critical time steps and Mamba-SSM modeling of long-range dependencies in complex multivariate interactions. While MAAT precision (97.48% vs. 97.14%) and recall (99.17% vs. 97.81%) both improve, rare red bars ($\Delta L_{i}<0$) align with transient noise where the baseline marginally outperforms. MAAT excels in Affiliation Recall (85.06% vs. 80.85%) and Range-based Precision (95.09% vs. 93.00%), ensuring precise localization of anomalies in noisy, high-dimensional sequences. The gated attention mechanism dynamically balances reconstruction fidelity and contextual enhancement, enabling robust performance in server monitoring applications where false positives and missed anomalies carry significant operational costs.
@@
7 Conclusions and future works
In this work, we introduced the Mamba Adaptive Anomaly Transformer (MAAT), which enhances unsupervised time series anomaly detection for practical applications like industrial monitoring and environmental sensing. MAAT effectively captures short- and long-term temporal dependencies by improving association discrepancy modeling with a new Sparse Attention mechanism and integrating a Mamba-Selective State Space Model (Mamba-SSM). Gated Attention mechanisms allow for a flexible combination of features from the original reconstruction and the Mamba-SSM output, striking a balance between Accuracy and context. Moreover, MAAT trains efficiently, making it suitable for resource-limited settings. Evaluation of benchmark datasets has revealed that MAAT significantly surpasses alternative methods, including the Anomaly Transformer and DCdetector, in both anomaly detection accuracy and generalization across various time series. It mitigates challenges such as sensitivity to short context windows and computational inefficiencies that have been prevalent in prior methodologies. Despite its efficiency in training, MAAT still has limitations. It is susceptible to hyperparameters, especially when balancing the reconstruction module with the Mamba-SSM pathway. Tuning these parameters is essential for optimizing performance across various datasets. Additionally, although we have designed our method to reduce the effects of noise and non-stationary conditions, its performance may decline in situations with very high noise or when abnormal patterns closely mimic normal ones. Finally, while MAAT shows reduced computational demands compared to some deep learning options, further enhancement of its inference speed is necessary for real-time applications. There is potential for future research in several areas. We could explore adaptive, data-driven hyperparameter tuning to stabilize the model and boost its performance in various conditions. Incorporating online learning or incremental updates could enhance MAAT functionality in dynamic environments where quick anomaly detection is crucial. Additionally, exploring hybrid models that combine the strengths of reconstruction-based approaches with contrastive learning may yield better results, especially in challenging situations characterized by significant non-stationarity or noise. By pursuing these avenues, future work can build on MAAT and advance unsupervised time series anomaly detection in real-world applications.
Acknowledgement

### v2

@@ 6.2 Reconstruction loss
where $L^{AT}_{i}$ and $L^{MAAT}_{i}$ represent losses for each batch $i$. The results are visualized in Figure 2, with green columns indicating lower loss for our method ($\Delta L_{i}>0$) and red columns showing a superior performance by AT ($\Delta L_{i}<0$). This illustration highlights the advantages of our approach during training.
MAAT outperforms AT on the SMD dataset with an F1 score of 92.30%, benefiting from effective reconstruction loss minimization (Figure 2(a)). The differential analysis mainly shows positive $\Delta L_{i}$, underscoring the effectiveness of Sparse Attention and Mamba-SSM in capturing critical time steps and long-range dependencies. Minor precision fluctuations do not compromise robustness, as evidenced by an Affiliation Recall of 93.51% and a Range-based Recall of 79%.
On the MSL dataset, MAAT also achieves lower reconstruction loss (Figure 2(a)), with a predominance of green bars in the $\Delta L_{i}$ plot, indicating its capacity to manage noisy patterns from Mars rover sensors. Despite slight precision trade-offs, MAAT shows significantly higher recall (96.03%) and robust Affiliation Recall (96.49%).
In the SMAP dataset, MAAT excels with an F1 score of 96.99%, consistently maintaining lower reconstruction loss (Figure 2(f)). The logarithmic difference reveals a dominance of green bars, highlighting MAAT’s proficiency in identifying subtle anomalies in satellite telemetry. Although there is a minor recall trade-off, precision gains enhance Affiliation Precision (49.34%).
Lastly, on the PSM dataset, MAAT achieves an F1 score of 98.32%, demonstrating its effectiveness with high-dimensional sensor data (Figure 2(c)). While facing occasional red bars in the $\Delta L_{i}$ plot, MAAT maintains superior Affiliation Recall (85.06%) and Range-based Precision (95.09%), ensuring accurate anomaly localization in noisy sequences.
7 Conclusions and future works
In this work, we introduced the Mamba Adaptive Anomaly Transformer (MAAT), which enhances unsupervised time series anomaly detection for practical applications like industrial monitoring and environmental sensing. MAAT effectively captures short- and long-term temporal dependencies by improving association discrepancy modeling with a new Sparse Attention mechanism and integrating a Mamba-Selective State Space Model (Mamba-SSM). Gated Attention mechanisms allow for a flexible combination of features from the original reconstruction and the Mamba-SSM output, striking a balance between Accuracy Accuracyext. Moreover, MAAT trains efficiently, making it suitable for resource-limited settings. Evaluation of benchmark datasets has revealed that MAAT significantly surpasses alternative methods, including the Anomaly Transformer and DCdetector, in both anomaly detection accuracy and generalization across various time series. It mitigates challenges such as sensitivity to short context windows and computational inefficiencies that have been prevalent in prior methodologies. Despite its efficiency in training, MAAT still has limitations. It is susceptible to hyperparameters, especially when balancing the reconstruction module with the Mamba-SSM pathway. Tuning these parameters is essential for optimizing performance across various datasets. Additionally, although we have designed our method to reduce the effects of noise and non-stationary conditions, its performance may decline in situations with very high noise or when abnormal patterns closely mimic normal ones. Finally, while MAAT shows reduced computational demands compared to some deep learning options, further enhancement of its inference speed is necessary for real-time applications. There is potential for future research in several areas. We could explore adaptive, data-driven hyperparameter tuning to stabilize the model and boost its performance in various conditions. Incorporating online learning or incremental updates could enhance MAAT functionality in dynamic environments where quick anomaly detection is crucial. Additionally, exploring hybrid models that combine the strengths of reconstruction-based approaches with contrastive learning may yield better results, especially in challenging situations characterized by significant non-stationarity or noise. By pursuing these avenues, future work can build on MAAT and advance unsupervised time series anomaly detection in real-world applications.
Acknowledgement
Appendix A Appendix: Hyperparameters for Model Training
Table 7 outlines the hyperparameters utilized for training the MAAT model across various datasets.The dimensions of the window, which set the length of the input sequence processed at each step, is set to 100 for most datasets. However, for SMAP, a slightly larger window of 105 is employed to better capture long-range dependencies. The batch size is tailored to the specific characteristics of each dataset and the computational constraints, with values ranging from 32 (for NIPS-TS-GECCO) to 256 (for MSL and SWaT).This range guarantees effective gradient updates while preserving training stability. The model dimension ($d_{\text{model}}$) remains consistently set at 512 for all datasets, defining the size of feature representations and ensuring consistency in learned embeddings. Additionally, the anomaly ratio, representing the proportion of data points identified as anomalies in each dataset, ranges from 0.5% to 1%, which affects the model’s responsiveness to infrequent events. These hyperparameters have been carefully chosen to enhance performance across various anomaly detection contexts.
Table 7: Hyperparameters for MAAT model across datasets
Dataset Window Size Batch Size $d_{\text{model}}$ Anomaly Ratio (%)
SMD 100 128 512 0.5
MSL 100 256 512 0.85
SMAP 105 128 512 0.5
PSM 100 128 512 1.0
SWaT 100 256 512 0.5
NIPS-TS-GECCO 100 32 512 0.5
NIPS-TS-SWAN 100 128 512 0.9
Appendix B Evalaution metrics
Evaluating anomaly detection in time series data requires metrics assepointwiset-wise performance and capturing anomalies’ temporal continuity. This work uses traditional and specialized point-based metrics to evaluate contiguous anomaly segments.
B.0.1 Point-Based Metrics
B.0.2 Affiliation Metrics: Aff-P and Aff-R
Affiliation Precision (Aff-P):
[…]
Affiliation Recall (Aff-R):
Measures the fraction of true anomalies detected:
[…]
B.0.3 Range-Based Metrics
These assess detection over entire anomaly segments.
Range-based Anomaly Recall (R_A_R):
[…]
Range-based Anomaly Precision (R_A_P):
[…]
B.0.4 Volume-Based Metrics: V_ROC and V_PR
These incorporate anomaly duration into evaluation.
Volume-based ROC (V_ROC):
Adjusts TPR and FPR by anomaly volume:
[…]
The V_ROC curve plots $\text{TPR}_{\text{vol}}$ vs. $\text{FPR}_{\text{vol}}$.
Volume-based Precision-Recall (V_PR):
Volume-weighted precision and recall are defined as follows:
In summary, while point-based metrics offer a basic assessment, affiliation, range-based, and volume-based metrics (11, 12, 13, 14) provide a comprehensive evaluation by accounting for temporal structure and anomaly duration [42, 43].
Appendix C Reconstruction loss
@@
Appendix D Implementation Details
The MAAT model was implemented using Python 3.8 and the PyTorch deep learning framework. PyTorch was chosen for its flexibility in defining custom architectures, efficient tensor operations, and strong support for GPU acceleration. The model was trained and evaluated on an NVIDIA RTX Titan GPU with 24GB VRAM, enabling efficient parallel computation for high-dimensional time-series data.
To ensure stable and reproducible training, experiments were conducted in a controlled environment with fixed random seeds. The training process leveraged mixed-precision computation to optimize memory usage and speed, particularly beneficial for large-scale datasets such as SMD and SWaT. The implementation also utilized PyTorch’s DataLoader for efficient batch processing, automatic differentiation via autograd, and hardware acceleration with torch.cuda.
The model was optimized using the Adam optimizer with weight decay regularization, and learning rate scheduling was applied to adapt to convergence dynamics. All experiments were run on a system equipped with Intel Xeon CPUs and 64GB RAM to handle large dataset preprocessing and training workloads efficiently.
Appendix E MAAT algorithm
This appendix contains details about the MAAT block algorithm used in the study.
1:Time series $x\in\mathbb{R}^{B\times L\times D}$
3:Reconstruction $x$, Anomaly scores: $series$, $prior$, $\sigma$
5:$x_{orig}\leftarrow x$triangleright Preserve initial input
6:for $i=1$ to $e\_layers$ do
7:     Sparse Attention Processing:
9:     Mamba Skip Path:
12:     $x_{skip}\leftarrow\text{LayerNorm}(x_{skip})$
13:     Adaptive Gating:
16:     State Update:
17:     $x_{orig}\leftarrow x$triangleright Update input for next layer
18:     Append $series$ to $series\_list$
19:     Append $prior$ to $prior\_list$
20:     Append $\sigma$ to $\sigma\_list$
21:end for
[…]
23:     $x\leftarrow\text{LayerNorm}(x)$
24:end if
25:return $x,series\_list,prior\_list,\sigma\_list$